OpenStudy (anonymous):

INTEGRATE 5x(sqrt(x+4))

OpenStudy (anonymous):

\[\int\limits_{}^{}5x(\sqrt{x+4})\]

zepdrix (zepdrix):

Oh ok now I remember what to do. This is one of those tricky ones. It's a simple u-sub

zepdrix (zepdrix):

\[\Large\rm u=x+4\]

OpenStudy (shiraz14):

Sub. u=sqrt(x+4)

zepdrix (zepdrix):

\[\Large\rm \int\limits5\color{orangered}{x}\sqrt{\color{royalblue}{x+4}}~\color{green}{dx}\]You will have 3 things you need to replace.

zepdrix (zepdrix):

ya or you can do that also^

OpenStudy (anonymous):

i did that. I ended up with\[\int\limits_{}^{}5(u-4)u ^{\frac{ 1 }{ 2 }}\]

zepdrix (zepdrix):

Ok good! Now distribute your u^(1/2) and integrate as you normally would! power rule to each term

OpenStudy (anonymous):

i ended up with\[2u ^{\frac{ 5 }{ 2 }}-\frac{ 40 }{ 3}u ^{\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

but then it differs by far from the wolfram answer, and its not good for our discussion here, because I really do not know what to say to the other guys now.

zepdrix (zepdrix):

No it's the same as wolfram's solution. You just have to tweak it.

zepdrix (zepdrix):

So if we factor a 2/3 and u^{3/2} out of each term we get,\[\Large\rm \frac{2}{3}u^{3/2}(3u-20)\]Lemme do that in my head really quick.. make sure I didn't goof up..

zepdrix (zepdrix):

Yah I think that's right. Then undoing our substitution gives us,\[\Large\rm \frac{2}{3}(x+4)^{3/2}(3(x+4)-20)\]Which simplifies a little bit further to match things up right?

zepdrix (zepdrix):

It can be tough to check your answers sometimes, wolfram and even textbooks tend to `over simplify` things.

OpenStudy (anonymous):

u r damn ryt man. have bn trying for hours!

zepdrix (zepdrix):

oh good we figured it out \c:/ heh

OpenStudy (anonymous):

cheers!